# DAFE: LLM-Based Evaluation Through Dynamic Arbitration for Free-Form Question-Answering

### v1

@@ Abstract
Evaluating Large Language Models (LLMs) free-form generated responses remains a challenge due to their diverse and open-ended nature. Traditional supervised signal-based automatic metrics fail to capture semantic equivalence or handle the variability of open-ended responses, while human evaluation, though reliable, is resource-intensive. Leveraging LLMs as evaluators offers a promising alternative due to their strong language understanding and instruction-following capabilities. Taking advantage of these capabilities, we propose the Dynamic Arbitration Framework for Evaluation (DAFE), which employs two primary LLM-as-judges and engages a third arbitrator only in cases of disagreements. This selective arbitration prioritizes evaluation reliability while reducing unnecessary computational demands compared to conventional majority voting. DAFE utilizes task-specific reference answers with dynamic arbitration to enhance judgment accuracy, resulting in significant improvements in evaluation metrics such as Macro F1 and Cohen’s Kappa. Through experiments, including a comprehensive human evaluation, we demonstrate DAFE’s ability to provide consistent, scalable, and resource-efficient assessments, establishing it as a robust framework for evaluating free-form model outputs.
@@ 1 Introduction
The rapid advancements in Large Language Models (LLMs) have propelled the field of natural language processing forward, yet their evaluation remains a challenge (Laskar et al., 2024). In particular, free-form model responses are difficult to evaluate because their correctness depends on understanding the broader context and underlying meaning (Si et al., 2021). Many benchmarks, such as MMLU (Hendrycks et al., 2021), often simplify evaluation by focusing on structured formats (e.g., multiple-choice questions) (Chen et al., 2024). Although effective for certain tasks, such methods rely on log probabilities assigned to predefined options, where the model selects the most likely answer, limiting the range of capabilities that can be assessed (Thakur et al., 2024). This structured approach fails to accommodate the complexity of free-form responses, where multiple valid answers exist (Chang et al., 2024). The rigid, predefined options in such evaluations not only limit the scope of assessment but also overlook the diversity of potential correct responses in free-form tasks (Li et al., 2023; Zhang et al., 2024).
Automatic metrics including lexical matching, n-gram, and neural-based have been widely adopted as scalable solutions for the evaluation of free-form model outputs. Lexical matching methods such as Exact Match (EM) evaluate model predictions by assessing strict lexical alignment between generated outputs and reference answers. However, EM fails to account for semantically equivalent variations in phrasing. For instance, despite their equivalence, EM treats “nuclear weapon” and “atomic bomb” as incorrect. Similarly, n-gram-based metrics (Papineni et al., 2002; Lin, 2004) primarily assess surface-level similarity and often fail to capture semantic equivalence, particularly when lexical or structural diversity conveys the same underlying meaning (Zhu et al., 2023; Chen et al., 2021; Zhang et al., 2020). Neural-based metrics like BERTScore (Zhang et al., 2020) address such limitations by leveraging contextual embeddings to evaluate semantic similarity. However, BERTScore depends on reference quality (Liu et al., 2024) and struggles with domain adaptation and length variations (Zhu et al., 2023). Furthermore, continuous score provider metrics are difficult to interpret (Xu et al., 2023). The limitations in automatic metrics become particularly evident when evaluating instruction-tuned chat models (Doostmohammadi et al., 2024), which tend to produce verbose and diverse responses (Saito et al., 2023; Wang et al., 2024b).
Contrary to automatic metrics, human evaluation provides a more transparent assessment (Chiang and Lee, 2023). However, despite being the “gold standard”, human evaluation is not without its limitations. LLMs’ growing complexity and scale have made recruiting and coordinating multiple human raters increasingly resource-intensive and time-consuming (Mañas et al., 2024). Furthermore, the reliability of human evaluation is additionally challenged by variations in rater expertise and inherent subjectivity that affect reproducibility (Clark et al., 2021; Chiang and Lee, 2023).
Recently, a paradigm shift has emerged where LLMs are utilized to judge the candidate model generations for given tasks (Zheng et al., 2024). This model-based method leverages the instruction-following capabilities of LLMs through evaluation prompts or, in some cases, fine-tuned versions of LLMs that are specifically optimized for evaluation. In this new line of work, research primarily focuses on pairwise comparison (Zheng et al., 2024; Wang et al., 2023; Vu et al., 2024), such as instructing an LLM to judge “which assistant response is better”, and single-answer scoring (Verga et al., 2024) like evaluating summarization task based on predefined criteria (e.g., likability, relevance, etc.) (Chiang and Lee, 2023; Hu et al., 2024; Liu et al., 2023; Chan et al., 2024; Chu et al., 2024).
Inspired by a recent study on self-correction where external feedback helps models identify and correct their mistakes (Gou et al., 2024a), we propose to guide LLM-as-a-judge with human-annotated task-specific reference answers in order to explore the potential of LLMs as an alternative to lexical matching (e.g., EM), neural-based (e.g., BERTScore), and human evaluation for automatic evaluation of free-form model responses. Unlike traditional metrics, an LLM judge can leverage its language understanding and instruction-following capabilities to recognize the correctness of open-ended generations.
We propose the Dynamic Arbitration Framework for Evaluation (DAFE), which employs LLM judges to evaluate free-form model responses. Using a single LLM as a judge, while simple, often leads to inconsistent evaluations, undermining trust in the results. On the other hand, the common practice of using large, universally capable models such as GPT-4 as evaluators makes the evaluation process both slow and costly (Jung et al., 2024; Adlakha et al., 2024; Verga et al., 2024), further limiting its broader applicability. Relying on multiple judges for every evaluation, though more reliable, exacerbates these computational challenges, making such approaches impractical at scale. DAFE offers a middle ground between these approaches by utilizing two complementary primary judges to perform the initial assessment. Only when these judges disagree, is a third independent arbitrator engaged to resolve the conflict. This selective arbitration ensures evaluation reliability and fairness while reducing computational overhead. Our experiments reveal that DAFE achieves significant improvements in metrics such as Macro F1 and Cohen’s kappa. Our key contributions include: a detailed analysis of limitations in conventional metrics for free-form QA, an evaluation of LLM judges with insights into their strengths and errors, a comprehensive human evaluation for benchmarking, and the introduction of DAFE—a scalable framework that improves reliability while minimizing the need for additional evaluators through selective arbitration.
Figure 1: Our proposed Dynamic Arbitration Framework for Evaluation (DAFE). Two primary judges, $J_{1}$ and $J_{2}$, first provide verdicts $V_{i_{1}}$ and $V_{i_{2}}$ for an instance $i$. If agree, that consensus $V_{i}$ is the final decision $D_{i}$. If disagree, a tiebreaker model $J_{t}$ independently produces a verdict $V_{t}$. The final decision $D_{i}$ is then determined via majority voting among $\{V_{i_{1}},V_{i_{2}},V_{t}\}$.
@@ 2 Methodology
This section briefly describes the key components of our proposed framework.
2.1 Candidate LLMs
A candidate LLM $\mathcal{C}_{\text{llm}}$ generates output $\bar{y}$ for the given input $x$. We first utilized candidate LLMs to obtain outputs for the given free-form question-answering tasks.
2.2 LLMs-as-a-Judge
A judge $\mathcal{J}_{\text{llm}}$ LLM delivers evaluation or verdict $V$ on candidate LLMs $\mathcal{C}_{\text{llm}}$ outputs $\bar{y}$. The $\mathcal{J}_{\text{llm}}$ evaluates output when prompted with $x$ (i.e., $x\rightarrow\mathcal{A}_{\text{llm}}$) and $\bar{y}$. We utilized the reference answer $r$ and prompted $P$ the $\mathcal{J}_{\text{llm}}$ as:
$P=\{x,\bar{y},r\}$
Utilizing $P$, $\mathcal{J}_{\text{llm}}$ performs the evaluation and delivers a decision as $V=J(P)$. The structure of this $V$ depends on the instructions provided in $P$. For instance, if a binary $V$ is required, $J$ assesses whether $\bar{y}$ is aligned with $r$ given the context $x$ and returns True if $\bar{y}$ is deemed correct, or False if it is not. The evaluation $P$ may vary from zero-shot, where $\mathcal{J}_{\text{llm}}$ receives no prior examples, to few-shot, which includes several related examples, or a chain of thought, encouraging $\mathcal{J}_{\text{llm}}$ to reason stepwise through the problem.
2.3 Dynamic Arbitration Framework for Evaluation (DAFE)
In traditional human evaluation settings, when two annotators disagree on a judgment, a third expert is often called upon to resolve the dispute. Drawing inspiration from this efficient human arbitration practice, we propose the Dynamic Arbitration Framework for Evaluation (DAFE). Rather than immediately employing a large powerful or a closed-source LLMs-as-a-judge, DAFE adopts a cost-efficient approach by beginning with two complementary open-source models as primary judges based on their past performance (Kenton et al., 2024). When these judges reach a consensus, no further evaluation is needed. Only in cases of disagreement is the more powerful LLM engaged as an arbitrator, whose decision then creates a majority verdict. This dynamic approach maintains evaluation quality while minimizing reliance on expensive models. The method also accounts for varying skill levels across different LLMs and tasks (Liang et al., 2024; Sun et al., 2024).
Formally, let $V_{i_{1}}$ and $V_{i_{2}}$ denote the verdicts from the two primary judges for the $i$-th evaluation instance. We define the agreement status $A_{i}$ as:
$A_{i}=\begin{cases}1&\text{if }V_{i_{1}}=V_{i_{2}},\\[6.0pt] 0&\text{otherwise}.\end{cases}$
If $A_{i}=1$, the final decision $D_{i}$ is simply $V_{i}$, the agreed-upon verdict of the primary judges. If $A_{i}=0$, a tiebreaker model provides an additional verdict $V_{t}$. The final decision $D_{i}$ is then obtained via majority voting among $\{V_{i_{1}},V_{i_{2}},V_{t}\}$. Formally:
$D_{i}=\begin{cases}V_{i}&\text{if }A_{i}=1,\\[6.0pt] \text{majority}(\{V_{i_{1}},V_{i_{2}},V_{t}\})&\text{if }A_{i}=0.\end{cases}$
The majority operation selects the verdict that appears at least twice among $\{V_{i_{1}},V_{i_{2}},V_{t}\}$. Since there are three votes, at least two must coincide for a majority.
@@ 3 Experiments
We utilize the following settings to examine the performance and reliability of individual LLM judges and DAFE.
3.1 Models
We select open and closed-source instruct models to serve as candidates and judges in our experiment. These include DeepSeek-V3 (DeepSeek-AI et al., 2025), Llama-3.1 70B [^1] https://huggingface.co/meta-llama/Meta-Llama-3.1-70B-Instruct (Meta AI, 2024), GPT-3.5-turbo (Brown et al., 2020), Mistral 7B [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), and Mixtral 8x7B [^3] https://huggingface.co/mistralai/Mixtral-8x7B-Instructv0.1 (Jiang et al., 2024). We also utilize GPT-4o (OpenAI et al., 2023) and DeepSeek-R1 (DeepSeek-AI et al., 2025) in our ablation experiments. To ensure the reproducibility of our experiments, we set the temperature to 0 for all models under study, as the performance of LLM-based evaluators has been shown to drop when temperature increases (Hada et al., 2024). For our proposed DAFE method, we utilized Mistral 7B and Llama 3.1 70B as primary judges with GPT-3.5-turbo as the tiebreaker. In addition, we experiment with other models as tiebreakers in our ablation experiments. In the rest of the paper, we refer both candidate and judge LLMs as: DeepSeek, Llama, GPT, Mistral, and Mixtral.
3.2 Datasets
We focus on free-form question-answering (QA) since it has widespread practical applications and the critical importance of truthfulness in this domain (Gou et al., 2024a; Evans et al., 2021). In our experiment, we utilize five free-form QA datasets: AmbigQA (Min et al., 2020), FreshQA (Vu et al., 2023), HotpotQA (Yang et al., 2018), Natural Questions (Kwiatkowski et al., 2019), and TriviaQA (Joshi et al., 2017). See Appendix A for details.
3.3 Prompts
We designed generalized (i.e., with minimum instructions) zero-shot prompts with role-playing (Kong et al., 2024) for both candidates and judges. Initially, we prompt candidate LLMs to elicit outputs for the given random samples associated with each dataset.
To evaluate the outputs of candidate LLMs, we prompt judge LLMs for binary verdicts (i.e., True or False) using $P=\{x,\bar{y},r\}$ and instructed to provide a brief explanation for their verdicts (see Appendix D for examples). Binary verdicts explicitly differentiate between correct and incorrect answers, minimize subjective interpretations, and simplify the evaluation process, thus facilitating automatic evaluation. In addition to three key prompt components (i.e., $x$, $\bar{y}$, $r$), we define the role of the judge LLMs as “You are a helpful assistant acting as an impartial judge.” to mitigate biases in judgments (Zheng et al., 2024). We chose not to use few-shot or chain-of-thought prompting strategies to keep the solution robust to a variety of tasks. Previous studies have also shown that in-context examples do not significantly improve the performance of model-based evaluators (Hada et al., 2024; Min et al., 2022).
3.4 Baselines
We establish the following baselines.
Exact Match (EM):
For our selected datasets and also free-form QA tasks, EM serves as a standard lexical matching metric to evaluate candidate LLM performance (Izacard and Grave, 2021; Lewis et al., 2020; Gou et al., 2024b). Due to the verbose nature of LLM-generated responses, we adapt EM to classify an answer as correct if any golden answer $r_{i}\in R$ appears within the generated response $\bar{y}$ (i.e., $r_{i}\subseteq\bar{y}$), rather than requiring complete strict string equality (i.e., $\bar{y}=r_{i}$).
BERTScore:
We use BERTScore (Zhang et al., 2020) which measures similarity by comparing contextualized word embeddings derived from a pre-trained BERT model. This enables the evaluation to focus on semantic correctness rather than exact lexical matches. As BERTScore is based on continuous values between -1 and 1, we set a threshold of $\tau=0.5$ to convert continuous similarity scores into binary 0 and 1. The purpose of this conversion is to allow direct comparison with other evaluation methods. For our implementation, we use the microsoft/deberta-xlarge-mnli [^4] https://huggingface.co/microsoft/deberta-xlarge-mnli model (He et al., 2021).
G-Eval:
In addition to automatic metrics, we also utilize G-Eval (Liu et al., 2023), a reference-free framework that uses GPT-4 to assess the quality of the generated text. In this setting, we modify the evaluation prompt by excluding the reference answer $r$ and directly prompted the evaluator model as $P=\{x,\bar{y}\}$ along with instructions.
Human Evaluation:
It remains the gold standard for assessing the outputs of candidate LLMs. We recruit three graduate students from our academic network, all specialized in natural language processing, to serve as annotators. We provide the input given to the candidate LLMs, reference answers, and candidate LLMs responses. This format, while similar, is distinct from the judge LLMs prompts which additionally require formatted decisions. We anonymize the origin of model responses to reduce potential bias linked to model familiarity or reputation. The annotators were asked to score the candidate LLMs outputs on a binary scale: ‘1’ for ‘True’ and ‘0’ for ‘False’ based on alignment with the reference answer and contextual relevance. For inter-rater reliability, we compute Fleiss’ Kappa ($\kappa$(Fleiss and Cohen, 1973) and percent agreement. See Appendix B for details.
@@ 4 Results
Figure 2 illustrates the raw performance of Llama obtained through various evaluators. Unlike lexical matching and neural-based metrics, each LLM-as-a-judge shows overall performance close to the human majority. The proposed DAFE method consistently achieves comparable or slightly better alignment with the human majority. Conventional metrics such as EM severely underestimate the candidate LLMs’ performance. Contrarily, BERTScore tends to overestimate the performance except in some cases such as when evaluating Llama on AmbigQA and NQ-Open (see Table 6 in Appendix C for additional results).
Figure 2: Raw accuracy of candidate Llama across free-form QA tasks using automatic metrics and model-based evaluation. The Human Majority (HM) serves as the ground truth for all evaluators.
4.1 Alignment with human evaluation
We calculate Cohen’s kappa (McHugh, 2012) to find the agreement between each evaluator and the human majority to obtain instance-level comparison. Overall, DAFE is almost perfectly aligned with human judgment than other evaluators (see Table 1). Similarly, individual LLM judges show more substantial to nearly perfect agreement with human judgments than EM and BERTScore.
Due to the high-class imbalance in TriviaQA, kappa scores can be misleadingly low despite high raw agreement - a known limitation called the “kappa paradox” (Cicchetti and Feinstein, 1990). Therefore, we treat the evaluation as a binary classification task where we consider each evaluator’s predictions against the human majority and report Macro-F1 scores which give equal weight to both classes regardless of their frequency in the selected random samples.
Table 1: Cohen’s Kappa scores displaying the agreement levels of various evaluators with human judgments across candidate models and tasks. Higher scores indicate better agreement with human judgments.

### v2

# CLEV: LLM-Based Evaluation Through Lightweight Efficient Voting for Free-Form Question-Answering

@@ Abstract
Evaluating free-form Question-Answering (QA) remains a challenge due to its diverse and open-ended nature. Traditional automatic metrics fail to capture semantic equivalence or accommodate the variability of open-ended responses. Leveraging Large Language Models (LLMs) as evaluators offers a promising alternative due to their strong language understanding and instruction-following capabilities. We propose the Consensus via Lightweight Efficient Voting (CLEV), which employs two primary LLMs as judges and engages a third judge only in cases of disagreement. This approach prioritizes evaluation reliability while reducing unnecessary computational demands. Through experiments, including human evaluation, we demonstrate CLEV’s ability to provide consistent, scalable, and resource-efficient assessments, establishing it as a robust framework for evaluating LLMs on free-form QA.
@@ 1 Introduction
One of the practical limitations in evaluating free‑form Question‑Answering (QA) is the lexical–semantic mismatch between Large Language Model (LLM) outputs and the reference answers. For the query “Who wrote 1984?”, a dataset may list simply “George Orwell”, while a helpful model replies: “It was penned by the British author Eric Arthur Blair.” Although both refer to the same person, there is no surface token overlap, so lexical-matching and even embedding-based metrics assign it an unduly low score. For instance, Exact Match (EM) requires strict lexical alignment (e.g., failing to equate “nuclear weapon” and “atomic bomb”) and ignores semantic equivalence (doostmohammadi2024reliable).
A reference-aware LLM-as-a-judge can instead reason over meaning, recognize that the candidate entails the gold fact, and deliver a verdict, thereby overcoming this lexical–semantic gap (10.5555/3666122.3668142; verga2024replacing). Existing studies using LLM as judges primarily focus on subjective pairwise comparison (wang2023large; vu2024autoraterstaminglarge) and single-answer scoring (chianglee2023large; hu2024llmbased; liuetal2023g). However, objective evaluation using LLM judges, particularly for free-form QA, has received limited attention. Furthermore, LLM-based judging also lies on a cost–quality spectrum. Querying a single judge is efficient but is less reliable due to the known limitations, such as prompt sensitivity, inconsistency, and bias (ye2024justiceprejudicequantifyingbiases), as no individual model captures the full diversity of reasoning styles, long-tail knowledge, and user values (feng2025llmdroolsmultillmcollaboration). To improve robustness, some studies employ fixed ensembles of three or more LLM judges and aggregate their decisions via majority voting (badshah-sajjad-2025-reference; verga2024replacing). While this improves reliability, it increases computational cost and latency (2024quantifyingcapabilitiesllmsscale), limiting scalability for large-scale evaluation (jung2024trustescalatellmjudges; adlakha2024evaluating).
To address these trade-offs, we propose the Consensus via Lightweight Efficient Voting (CLEV) that balances the reliability and efficiency of using LLMs as judges for free-form QA. CLEV employs two primary judges for initial assessments and invokes a third judge only when disagreements occur. By minimizing redundant calls in the fixed three-judge majority vote setting, CLEV reduces computational overhead by roughly 80 to 95% (varying by task) while achieving substantial to perfect agreement. Our key contributions include: 1) establishing a principled LLM-as-a-judge evaluation setup for free-form QA that moves beyond token-level string matching metrics, 2) proposing CLEV, a lightweight consensus-based evaluation method that preserves the reliability of multi-judge voting while significantly reducing computational cost, 3) empirically validating CLEV across diverse QA datasets and multiple state-of-the-art LLMs, and 4) systematic analysis of failure cases.
Figure 1: Our proposed Consensus via Lightweight Efficient Voting.
@@ 2 Methodology
This section outlines the core components of our evaluation framework. Given a question and a candidate model’s answer, the evaluation task is to determine whether the answer is factually correct with respect to a reference. Because free-form QA requires an objective correctness decision rather than scalar scoring, we prompt LLM judges to return a binary verdict along with a brief rationale explaining their judgment. The verdict serves as the evaluation signal, while the rationale provides interpretability and supports later analysis.
2.1 Consensus via Lightweight Efficient Voting (CLEV)
In traditional human evaluation settings, when two annotators disagree on a judgment, a third expert is often called upon to resolve the dispute. Drawing inspiration from this efficient practice, we propose CLEV. As illustrated in Figure 1, rather than immediately employing three LLMs, CLEV adopts an efficient approach by beginning with two models as primary judges. When these judges reach a consensus, no further evaluation is needed. Only in cases of disagreement, the third LLM is engaged, whose decision then creates a majority verdict.
2.2 Judges inclusion and exclusion criteria
To systematically select suitable judges for CLEV, we evaluate various LLMs (see Figure 2) using 100 random instances from HotpotQA. For each model, we compare binary verdicts against human annotations and compute Cohen’s Kappa ($\kappa$) and Macro F1. We interpret $\kappa$ following the commonly used guideline where values between 0.610.610.800.80 indicate substantial agreement, and values above $0.80$ indicate near-perfect agreement (mchugh2012interrater). However, since $\kappa$ is known to be sensitive to class imbalance (cicchetti1990high), we together consider Macro F1 to ensure balanced evaluation across both classes:
$\text{status}=\begin{cases}\text{primary judges},&\kappa\geq 0.6\ \text{and}\ \text{F1}\geq 0.85,\\[4.0pt] \text{third judge},&\kappa\geq 0.8\ \text{and}\ \text{F1}\geq 0.9,\\[4.0pt] \text{excluded},&\text{otherwise}.\end{cases}$
Figure 2: Judges selection based on the defined criteria in Section $2.$
@@ 3 Experiments
We utilize the following settings to evaluate CLEV:
Models
Our candidate models includes Llama-3.1 70B (MetaLlama3_2024), GPT-3.5-turbo (NEURIPS2020_1457c0d6), Mistral 7B (jiang2023mistral), and Mixtral 8x7B (jiang2024mixtralexperts). For judges, Mistral‑7B met our inclusion criteria ($\kappa\geq 0.6$, $\mathrm{F1}\geq 0.85$) while offering low cost and fast inference. Llama‑3.1‑70B, though more expensive, provided substantial agreement. We therefore use both as primary judges to capture a range of capabilities (feng2025llmdroolsmultillmcollaboration; liangetal2024encouraging; sun2024skillaggregation). GPT‑3.5‑turbo serves as the third judge because it exceeds the required thresholds ($\kappa\geq 0.8$, $\mathrm{F1}\geq 0.9$). All models are run with temperature 0 to ensure reproducibility, as higher temperatures degrade evaluator reliability (hada2024llmbasedevaluators).
Datasets
We utilize four datasets: AmbigQA (minetal2020ambigqa), HotpotQA (yang2018hotpotqa), Natural Questions (kwiatkowskietal2019), TriviaQA (joshi2017triviaqa). We only utilize FreshQA to evaluate the judge’s ability to detect outdated information. See Appendix A for details.
Prompts
We design minimal zero-shot role-playing prompts (kong2024groleplay) for both candidate and judge LLMs. The judge LLM is instructed to return a binary verdict with a brief explanation, which reduces subjectivity and simplifies automatic evaluation (see Appendix F).
Baselines
We compare individual LLM judges and CLEV against EM, BERTScore (ZhangKWWA20) computed with microsoft/deberta-xlarge-mnli and thresholded at $\tau=0.5$ to yield binary decisions and majority voting, which always invokes a fixed setup of LLM judges and returns the model verdict (see Appendix B).
Human evaluation
We invite three volunteer graduate students from our lab to act as annotators. We provided the questions, reference answers, and candidate LLM responses, without information about model identity, to avoid bias. Each response is scored on a binary scale based on correctness and relevance (see Appendix C).
Evaluation metrics
We compute Fleiss’ Kappa ($\kappa$(fleiss1973equivalence) and percent agreement to measure inter-rater reliability among human annotators. To compare evaluators with humans, we calculate Cohen’s Kappa (mchugh2012interrater) against the human majority on an instance level. Because of class imbalance, kappa can appear low despite high agreement (cicchetti1990high). Therefore, we also frame the task as binary classification and report Macro-F1 scores.
@@ 4 Results
In this section, we briefly report the results and refer the readers to Appendix E for detailed results.
Evaluators (Cohen’s $\kappa$ [Macro-F1] vs Human Majority)
Candid. Tasks EM BERTScore Llama GPT Mistral MV CLEV Disagr. (%)
Llama AmbigQA 0.52 [0.74] 0.28 [0.64] 0.89 [0.94] 0.84 [0.92] 0.86 [0.93] 0.91 [0.96] 0.91 [0.96] 10.0
HotpotQA 0.58 [0.78] 0.50 [0.75] 0.88 [0.94] 0.90 [0.95] 0.83 [0.92] 0.95 [0.98] 0.95 [0.98] 13.0
NQ-Open 0.38 [0.65] 0.44 [0.72] 0.83 [0.92] 0.79 [0.90] 0.74 [0.87] 0.93 [0.96] 0.92 [0.96] 18.0
TriviaQA 0.28 [0.61] 0.56 [0.78] 0.55 [0.77] 0.44 [0.72] 0.30 [0.64] 0.68 [0.84] 0.68 [0.84] 17.0
GPT AmbigQA 0.56 [0.79] 0.25 [0.62] 0.94 [0.97] 0.90 [0.95] 0.85 [0.93] 0.97 [0.98] 0.96 [0.98] 7.0
HotpotQA 0.60 [0.79] 0.30 [0.62] 0.95 [0.98] 0.97 [0.99] 0.93 [0.97] 0.99 [0.99] 0.98 [0.99] 5.7
NQ-Open 0.45 [0.70] 0.22 [0.61] 0.88 [0.94] 0.82 [0.91] 0.83 [0.91] 0.96 [0.98] 0.95 [0.98] 13.0
TriviaQA 0.34 [0.65] 0.36 [0.68] 0.65 [0.82] 0.40 [0.70] 0.47 [0.73] 0.81 [0.90] 0.77 [0.89] 15.7
Mixtral AmbigQA 0.55 [0.76] 0.34 [0.67] 0.90 [0.95] 0.78 [0.89] 0.89 [0.94] 0.98 [0.99] 0.95 [0.98] 9.0
HotpotQA 0.55 [0.76] 0.35 [0.66] 0.94 [0.97] 0.93 [0.97] 0.94 [0.97] 0.97 [0.99] 0.97 [0.99] 4.7
NQ-Open 0.37 [0.65] 0.30 [0.65] 0.88 [0.94] 0.73 [0.86] 0.82 [0.91] 0.97 [0.98] 0.91 [0.96] 13.0
TriviaQA 0.32 [0.63] 0.39 [0.70] 0.63 [0.81] 0.61 [0.80] 0.44 [0.72] 0.90 [0.95] 0.76 [0.88] 17.0
Mistral AmbigQA 0.60 [0.79] 0.25 [0.62] 0.89 [0.95] 0.89 [0.95] 0.86 [0.93] 0.95 [0.98] 0.95 [0.98] 11.7
HotpotQA 0.61 [0.80] 0.38 [0.67] 0.94 [0.97] 0.90 [0.95] 0.94 [0.97] 0.96 [0.98] 0.95 [0.98] 6.0
NQ-Open 0.48 [0.73] 0.29 [0.64] 0.85 [0.93] 0.84 [0.92] 0.84 [0.92] 0.95 [0.98] 0.95 [0.98] 14.7
TriviaQA 0.47 [0.72] 0.24 [0.61] 0.76 [0.88] 0.73 [0.86] 0.47 [0.74] 0.87 [0.94] 0.85 [0.93] 20.3
Table 1: Cohen’s Kappa and [Macro-F1] scores showing the agreement of individual evaluators, Majority Vote (MV), and CLEV with human majority across models and tasks. Disagr. (%) indicates the percentage of disagreements in the majority voting (Llama+GPT+Mistral), i.e., when the third judge is called in CLEV.
Correlation with human evaluation
As depicted by consistently high Cohen’s kappa in Table 1, CLEV maintains strong agreement with human evaluation. This represents an improvement over individual model performance, where individual judges generally showed varying levels of agreement with human evaluation.
LLM-based evaluators demonstrate strong abilities in recognizing semantic variations while maintaining the core meaning, especially when assessing responses that use different terminology or structural approaches to convey the same information. For instance, evaluators correctly identified that “Salma Hayek” and “Salma Hayek Pinault” refer to the same individual, acknowledging the semantic equivalence despite differences in phrasing. Similarly, when assessing responses that use different terms for the same entity, such as recognizing “Nick Fury, Agent of S.H.I.E.L.D.” as part of the broader “Marvel” universe, the evaluators effectively maintain the core meaning and contextual relevance. Their explanations show systematic assessment patterns that combine multiple evaluation criteria (e.g., factual accuracy and contextual relevance).
CLEV reducing third-judge calls by 80–95% while maintaining the performance across tasks. As depicted in Table 1, even on TriviaQA, where disagreement is high, CLEV scores remain in substantial to perfect agreement.
EM underestimates and BERTScore overestimates. Interestingly, EM typically accomplishes better correlation with human evaluation on the instance-level in Table 1 than neural-based BERTScore. EM’s strict nature leads to lower overall performance, but its high precision ensures that when it identifies a match, it strongly aligns with human annotations. In contrast, BERTScore takes a more lenient approach to semantic matching. As a result, it shows more false positives, consequently reducing instance-level agreement with human judgments.
4.1 Impact of Selective Third Judge
A core benefit of CLEV is that it avoids unnecessary third-judge evaluations by only invoking a judge when the two primary judges disagree. Table 2 summarizes the frequency of such escalation events across five datasets and five candidate LLMs. Across all 7,500 evaluated instances (300 examples $\times$ 5 models $\times$ 5 tasks), only 17.6% required a third judge. This reduction is substantial when compared to the fixed three-judge majority voting setup, which always incurs 100% of third-judge calls. In other words, CLEV reduces evaluation overhead by approximately 80–95% while preserving agreement levels that match majority vote outcomes (see Table 1). Tasks with stable factual grounding (e.g., HotpotQA) trigger few escalations, whereas more temporally dynamic datasets like FreshQA show higher disagreement rates due to shifts in real-world facts.
Tasks Third judge calls (of 1,500) Rate (%)
AmbigQA 190 12.7
FreshQA 543 36.2
HotpotQA 120 8.0
NQ-Open 212 14.1
TriviaQA 253 16.9
Total 1,318 / 7,500 17.6
Table 2: Aggregated third-judge usage under CLEV across tasks. Each task has 1,500 evaluation instances per candidate model (300 examples $\times$ 5 candidate LLMs), and the reported counts sum over all five models  [^2] We also included DeepSeek as our fifth model in the Appendix. See Table 8.. Lower values indicate fewer disagreement-triggered escalations.
4.2 Error analysis
We randomly sampled 100 error cases (50 false positives and 50 false negatives) from each evaluator to study their behavior. Given EM had 11 false positives, we included all of them in our analysis.
LLMs are prone to hallucination in justification where they fabricate reasoning to support their evaluations, and produce detailed but incorrect explanations. In LLM judges, false positives and negatives often result from overlooking critical distinctions between candidate LLM outputs and failing to account for the specificity required by the reference answer. This pattern is particularly noticeable in Mistral 7B, where the model disregards the ground truth and provides evaluations influenced by unknown factors. For example, when evaluating candidate GPT-3.5’s response “The foreign minister of Germany who signed the Treaty of Versailles was Hermann Müller.” which is correct according to the reference answer “Hermann Müller” and human evaluation, Mistral 7B as-a-judge incorrectly marked this response as false and fabricated reasoning in support of its decision.
Specificity bias occur in some judges. This approach shifts the evaluation towards false negatives by missing semantically similar but structurally different answers. We found many cases when such evaluators failed to account for valid variations in phrasing or granularity, focusing instead on rigid adherence to the reference answer. Compounding these issues are reasoning errors within the evaluators’ own explanations, which often contain overconfident assertions.
We found several temporal limitations in LLM-based evaluators. Despite using mostly older datasets with up-to-date models, we observed failures on recent events or evolving contexts. The FreshQA dataset (vu2023freshllms), being recent, highlights these temporal limitations. As shown in Table 3, LLM judges deviate more from human judgment on FreshQA compared to older datasets like HotpotQA. In dynamic or time-sensitive contexts, they often hallucinate by marking incorrect responses as True. For instance, when asked: “On what date did the Patriots last play the Miami Dolphins?”, the evaluator accepted an outdated answer, “January 1, 2023,” as correct, ignoring the reference “November 24, 2024,” and justifying its mistake. Because multiple judges rely on this same outdated parametric knowledge, FreshQA exhibits substantially higher disagreement between judges, who often converge on answers that were once correct but no longer reflect the current ground-truth.